# Lan-grasp: Using Large Language Models for Semantic Object Grasping

### v2

@@ 4 Experimental Evaluation
In this section, we present the details of our experiments and results. Our goal is to demonstrate that our method proposes to grasp object parts that are preferred by humans on a variety of objects. We argue that humans generally choose grasps that enable correct tool usage and ensure safety and that a robot retains these desirable qualities by executing similar grasps. To that end, we first collect a data set of typical household objects. Next, we apply our approach to these objects and execute the grasping on a real robot. Finally, we show that our grasping strategy is similar to human preferences obtained through a survey and that our approach outperforms two baselines based on this similarity metric. In the following, we describe our data set, provide details on the baseline approaches and the performed experiments, and discuss the results. In subsection 4.5 we present an extension to the main algorithm that reasons about the feasibility of a grasp in complex scenarios. Finally, in subsection 4.6 we perform an ablation study on several components of the pipeline.
@@ 4.1 Dataset
We collect a data set containing $22$ different objects commonly found in household environments. We chose these objects to cover a wide range of situations where semantic knowledge is required for proper grasping. Our first objective was to showcase grasping on functional objects like tools or kitchen supplies, e.g., shovel, hand brush, and knife. Further, we included delicate objects that might be damaged with an improper grasp, for instance, rose, cupcake, and ice cream. For other objects, a wrong grasp can cause a dangerous situation, e.g., candle. Finally, we include objects where an improper grasp might not necessarily be harmful but is rather unnatural to a human observer, for instance, doll, bag, and wine glass. The objects in the data set are shown in Figure 3, Figure 4, and Figure 5.
@@ 4.2 Experimental Setup and Baselines
For $11$ objects from our data set, listed in Table 1, we perform real-world experiments using the Human Support Robot (HSR) [46]. we first scan each object and than apply GraspIt! to the resulting 3D model as detailed in subsection 3.2. From the top-$20$ grasp proposals we randomly pick one and execute it using the proprietary HSR motion planner.
Our first baseline is the plain GraspIt! simulator. Here we use the same 3D models as for our approach but do not restrict grasping to the object part selected by the language model. For each object, we evaluate the top-$20$ grasp proposals and carry two of them out on the HSR, as shown in Figure 3.
The second baseline is GraspGPT [9], a recent approach to task-oriented grasping. This method requires as input an object point cloud and a natural language prompt describing the object, the object class, and the task. We generate the point clouds from the object meshes reconstructed as above and use an object-specific activity as the task label. Again, we retrieve $20$ grasps per object but do not carry them out on the robot.
@@ 4.3 Qualitative Results
In this section, we present and discuss the grasping results of Lan-grasp and the GraspIt!-baseline. The grasps executed on the robot are shown in Figure 3. For the rest of the objects, the grasping area suggested by our method is presented in Figure 5.
@@
Figure 3: The performed grasps by the HSR robot: Each column presents the grasps for one object. The first two rows for each object, show the grasps generated without semantic knowledge about the objects, while the third and fourth rows show the grasps generated by Lan-grasp.
Figure 4: The performed grasps by the HSR robot: Each column presents the grasps for one object. The first two rows for each object, show the grasps generated without semantic knowledge about the objects, while the third and fourth rows show the grasps generated by Lan-grasp.
@@
Object Preferred Part GraspIt! GraspGPT Lan-grasp
doll torso 92.1% 0.28 0.48 0.92
ice cream cone 100.0% 0.05 0.40 1.00
candle base 93.1% 0.22 0.57 0.93
flowers in the vase vase 93.2% 0.32 0.73 0.93
bag handle 91.1% 0.79 0.69 0.91
plant pot 94.3% 0.16 0.56 0.94
hand brush handle 95.4% 0.65 0.95 0.95
toilet brush handle 97.6% 0.42 0.52 0.98
cactus pot 98.8% 0.26 0.99 0.99
cupcake wrapper 100.0% 0.10 0.40 1.00
cup on a saucer saucer 81.2% 0.24 0.59 0.81
plate of cake plate 98.8% 0.11 0.51 0.99
mug handle 77.1% 0.28 0.73 0.77
saucepan handle 94.3% 0.36 0.94 0.94
broom handle 97.6% 0.42 0.98 0.98
@@
Table 1: Similarity of grasping area preferences compared to a human user. The left half of the table lists the objects and the object part the majority of survey participants suggested for grasping, with the corresponding percentage of users. The right half of the table shows the similarity scores per object for the two baselines and the proposed method.
@@ 4.4 Quantitative Results
To support the claim that our approach proposes grasps similar to human preferences, we designed a questionnaire on grasping choices. A group of $83$ participants were presented with images of all objects used in the experiments and were asked where they would grasp them. For each object, the participants could choose between two parts marked by bounding boxes in the image. The survey results are summarized in Table 1. Per object, we state the preferred part and the percentage of participants that selected it.
For the proposed approach and the baselines, we want to evaluate how similar the generated grasps are compared to the ones suggested by human users. Given that an object is segmented into parts $a$ and $b$, let $p_{a}\in[0,1]$ be the empirical probability that a method grasps at part $a$ and $p_{b}=1-p_{a}$ that part $b$ is grasped. Further, let $p_{a}^{h}$ be the human grasping frequency at $a$ according to the survey results and $p_{a}^{x}$ the corresponding frequency produced by one of the considered methods. To compute $p_{a}^{x}$ for the baselines, we obtained $20$ grasp proposals from each algorithm and counted the grasps falling into region $a$. Lan-grasp restricts the grasps to the object part selected by the LLM, which in our experiments robustly proposed the same part for a given object. Thus, the values of $p_{a}^{x}$ were here either $1$ or 00. Finally, we computed a per-object similarity score for each method $x$ as $sim_{x}=1-|p_{a}^{h}-p_{a}^{x}|$. These scores are shown in Table 1 along with the average similarity scores over all objects.
Our method consistently outperforms the baselines on the similarity score and ties only in four cases with GraspGPT. The average similarity score of Lan-grasp is considerably higher with the value of $0.94$ compared to $0.31$ achieved by GraspIt! and $0.67$ achieved by GraspGPT. We further note that in all cases, the object part choice of Lan-grasp coincides with the majority vote of the survey participants. The low score of GraspIt! is not surprising since it only considers geometric and not semantic aspects of the object. Thus, whether the object is grasped in a particular region is pure chance, and we expect the similarity score to be closer to $0.5$ for a larger data set. GraspGPT exhibits a better performance compared to GraspIt! due to leveraging semantic concepts and LLMs. However, it was tuned on a data set mostly containing tools and house supplies and thus does not perform well on objects like a doll or an ice cream.
@@ 4.5 Grasp Feasibility Feedback
So far, we assumed that the grasp suggested by the LLM is feasible, however, this is not always the case. For instance, the referenced object part might not be visible in the image, broken, or occluded by other objects. Further, the execution of the grasp could lead to undesirable outcomes. In order to mitigate this issue, we propose a feedback loop consisting of a VLM and an LLM that communicate with each other to find a feasible grasp. The approach and the used prompts are shown in Figure 6. As in our original pipeline, the LLM first suggests an object part based on the object label. Next, the VLM analyzes the image and evaluates the feasibility of the grasping part. Crucially, we follow the idea of Chain-of-Thought (CoT) prompting [47] and include the sentence "Start your answer with lets analyze the image." into the prompt. If feasible, we proceed as before. If not, the LLM receives a list of all object parts rejected so far and is asked to propose another grasp. This loop repeats until either a feasible grasp part is found or there are no more suitable parts of the object left for grasping. In our implementation, the roles of LLM and VLM are carried out by the same model (GPT-4o).
To evaluate this approach we gathered a dataset of $18$ challenging scenarios, shown in Figure 8 and Figure 7, where the initial grasp suggestion is not feasible and requires reasoning to find the right grasp part. We defined the ground truth manually to evaluate the results. We ran the algorithm $5$ times to obtain an average success rate of $91.14\%$ over all scenes. The qualitative results are shown for part of the dataset in Figure 7. Our algorithm can take into account different criteria, e.g., the risk due to the proximity of the lit candle, potential mess due to the overfilled cup, or occlusions by nearby objects. We further analyze the effects of VLM choice and CoT in the following section.
@@ 4.6 Ablation Study
In this section, we evaluate the influence of different algorithm components on the above results, specifically the choice of the LLM and VLM and the effect of CoT on the feasibility feedback. First, we consider the pipeline without feedback as described in subsection 3.1. Here we consider GPT-3.5-turbo and GPT-4 as LLMs. We further investigate whether replacing the LLM with a VLM improves the performance. To that end, we deploy GPT-4o, GPT-4o-mini, and the open-source LLaVA-1.5 7B VLM and provide them with an image of the object to grasp.
Here, we run our pipeline only until detecting the grasping area in the image, without executing the grasp nor generating a grasping pose. Therefore, we use different metrics than in subsection 4.4. First, we count the exact matches between the LLM-generated and GT object part labels and report the average success rate. Second, we compute the Intersection-over-Union (IoU) for the proposed grasping regions and the ground truth. All algorithm versions are evaluated on data from subsection 4.1 and the results are summarized in Table 2.
Both text-only GPT versions perform equally with an $81.8\%$ success rate and an IoU score of around $0.63$. The VLM variants perform slightly better with a success rate of $86.3\%$. We note that there is no difference between the flagship GPT-4o and the downsized GPT-4o-mini model. LLaVA performed significantly worse with $45.5\%$ success rate and $0.38$ IoU score. Analyzing the object parts suggested by LLaVA showed that the model was correct for objects possessing a handle, for other objects the answer was either wrong or referred to generic image locations, e.g., ’bottom’ or ’top’. A straightforward explanation could be simply the smaller model size. However, another reason might be that we used the same prompt for LLaVA as for the GPT models, and better results could be achieved with further prompt engineering specifically targeting LLaVA.
method text only text and image
@@
Table 2: Ablation of LLMs and VLMs in the main Lan-grasp pipeline.
For the feasibility feedback algorithm, we compare GPT-4o and GPT-4o-mini. Further, we experiment with two prompt variants, the first with CoT (zero-shot-CoT) and the second without (zero-shot). In the latter, we omit the sentence "Start your answer with lets analyze the image." from the prompt. The experiments were performed in the same fashion as in subsection 4.5 and the results are reported in Table 3. First, we consider the CoT variants. With $62.75\%$ success rate GPT-4o-mini performed significantly worse than the larger GPT-4o ($91.14\%$), which indicates that model size is an important factor for complex reasoning. Without CoT the performance dropped to $65.33\%$ for GPT-4o and $52.78\%$ for GPT-4o-mini. That result demonstrates that our algorithm, in fact, benefits from the CoT approach.
@@
success rate 91.14% 65.33% 62.75% 52.78%

### v3

@@ 4 Experimental Evaluation
In this section, we present the details of our experiments and results. Our goal is to demonstrate that our method proposes to grasp object parts that are preferred by humans across a variety of objects. We argue that humans generally choose grasps that enable correct tool usage and ensure safety, and that a robot retains these desirable qualities by executing similar grasps. To that end, we first collect a dataset of typical household objects. Next, we apply our approach to these objects and execute the grasping on a real robot. We show that our grasping strategy is similar to human preferences obtained through a survey and that our approach outperforms two baselines based on this similarity metric. Finally, we evaluate our proposed placement method on a range of different objects.
In the following sections, we describe our dataset, provide details on the baseline approaches and the experiments performed, and discuss the results. In Section˜4.5, we present an extension to the main algorithm that reasons about the feasibility of a grasp in complex scenarios. In Section˜4.6, we perform an ablation study on several components of the pipeline. Finally, in Section˜4.7 we present the setup details, results, and discussion for the object placement approach.
@@ 4.1 Dataset
We collect a dataset containing $22$ different objects commonly found in household environments. We chose these objects to cover a wide range of situations where semantic knowledge is required for proper grasping. Our first objective was to showcase the grasping of functional objects like tools or kitchen supplies, e.g., shovel, hand brush, and knife. Furthermore, we included delicate objects that might be damaged with an improper grasp, such as rose, cupcake, and ice cream. For other objects, a wrong grasp can cause a dangerous situation, e.g., candle. Finally, we included objects where an improper grasp might not necessarily be harmful but is rather unnatural to a human observer, for instance, doll, bag, and wine glass. The objects in the data set are shown in Figs.˜4, 5 and 6.
@@ 4.2 Experimental Setup and Baselines
For $11$ objects from our dataset, listed in Table˜1, we conduct real-world experiments using the Human Support Robot (HSR) (Yamamoto et al., 2019). We first scan each object and then apply GraspIt! to the resulting 3D model as detailed in Section˜3.2. From the top-$20$ grasp proposals, we randomly pick one and execute it using the proprietary HSR motion planner.
Our first baseline is the plain GraspIt! simulator. Here we use the same 3D models as for our approach but do not restrict grasping to the object part selected by the language model. For each object, we evaluate the top-$20$ grasp proposals and carry out two of them on the HSR, as shown in Fig.˜4.
The second baseline is GraspGPT (Tang et al., 2023a), a recent approach to task-oriented grasping. This method requires as input an object point cloud and a natural language prompt describing the object, the object class, and the task. We generate the point clouds from the object meshes reconstructed as above and use an object-specific activity as the task label. Again, we retrieve $20$ grasps per object but do not carry them out on the robot.
@@ 4.3 Qualitative Results
In this section, we present and discuss the grasping results of Lan-grasp and the GraspIt!-baseline. The grasps executed on the robot are shown in Figs.˜4 and 5. For the rest of the objects, the grasping area suggested by our method is presented in Fig.˜6.
@@
Figure 4: The grasps performed by the HSR robot: Each column presents the grasps for one object. The first two rows for each object show the grasps generated without semantic knowledge about the objects, while the third and fourth rows show the grasps generated by Lan-grasp.
Figure 5: The grasps performed by the HSR robot: Each column presents the grasps for one object. The first two rows for each object, show the grasps generated without semantic knowledge about the objects, while the third and fourth rows show the grasps generated by Lan-grasp.
@@
Table 1: Similarity of grasping area preferences compared to a human user. The left half of the table lists the objects and the object part the majority of survey participants suggested for grasping, with the corresponding percentage of users. The right half shows similarity scores for the two baselines and the proposed method.
Object Preferred Part (%) GraspIt! GraspGPT Lan-grasp
doll torso 92.1 0.28 0.48 0.92
ice cream cone 100.0 0.05 0.40 1.00
candle base 93.1 0.22 0.57 0.93
flowers in the vase vase 93.2 0.32 0.73 0.93
bag handle 91.1 0.79 0.69 0.91
plant pot 94.3 0.16 0.56 0.94
hand brush handle 95.4 0.65 0.95 0.95
toilet brush handle 97.6 0.42 0.52 0.98
cactus pot 98.8 0.26 0.99 0.99
cupcake wrapper 100.0 0.10 0.40 1.00
cup on a saucer saucer 81.2 0.24 0.59 0.81
plate of cake plate 98.8 0.11 0.51 0.99
mug handle 77.1 0.28 0.73 0.77
saucepan handle 94.3 0.36 0.94 0.94
broom handle 97.6 0.42 0.98 0.98
@@ 4.4 Quantitative Results
To support the claim that our approach proposes grasps similar to human preferences, we designed a questionnaire on grasping choices. A group of $83$ participants were presented with images of all objects used in the experiments and asked where they would grasp them. For each object, the participants could choose between two parts marked by bounding boxes in the image. The survey results are summarized in Table˜1. For each object, we state the preferred part and the percentage of participants who selected it.
For the proposed approach and the baselines, we aimed to evaluate how similar the generated grasps are to those suggested by human users. Given that an object is segmented into parts $a$ and $b$, let $p_{a}\in[0,1]$ be the empirical probability that a method grasps at part $a$ and $p_{b}=1-p_{a}$ that part $b$ is grasped. Further, let $p_{a}^{h}$ be the human grasping frequency at $a$ according to the survey results and $p_{a}^{x}$ the corresponding frequency produced by one of the considered methods. To compute $p_{a}^{x}$ for the baselines, we obtained $20$ grasp proposals from each algorithm and counted the grasps falling into region $a$. Lan-grasp restricts the grasps to the object part selected by the LLM, which, in our experiments, robustly proposed the same part for a given object. Thus, the values of $p_{a}^{x}$ were either $1$ or 0. Finally, we computed a per-object similarity score for each method $x$ as $sim_{x}=1-|p_{a}^{h}-p_{a}^{x}|$. These scores are shown in Table˜1 along with the average similarity scores over all objects.
Our method consistently outperforms the baselines on the similarity score and ties only in four cases with GraspGPT. The average similarity score of Lan-grasp is considerably higher with the value of $0.94$ compared to $0.31$ achieved by GraspIt! and $0.67$ achieved by GraspGPT. We further note that in all cases, the object part choice of Lan-grasp coincides with the majority vote of the survey participants. The low score of GraspIt! is not surprising since it only considers geometric and not semantic aspects of the object. Thus, whether the object is grasped in a particular region is purely a matter of chance, and we expect the similarity score to be closer to $0.5$ for a larger dataset. GraspGPT exhibits a better performance compared to GraspIt! due to leveraging semantic concepts and LLMs. However, it was tuned on a dataset mostly containing tools and house supplies and thus does not perform well on objects like a doll or an ice cream.
@@ 4.5 Grasp Feasibility Feedback
So far, we have assumed that the grasp suggested by the LLM is feasible. However, this is not always the case. For instance, the referenced object part may not be visible in the image, be broken, or be occluded by other objects. Further, the execution of the grasp could lead to undesirable outcomes. In order to mitigate this issue, we propose a feedback loop consisting of a VLM and an LLM that communicate with each other to find a feasible grasp. The approach and the used prompts are shown in Fig.˜7. As in our original pipeline, the LLM first suggests an object part based on the object label. Next, the VLM analyzes the image and evaluates the feasibility of the grasping part. Crucially, we follow the idea of Chain-of-Thought (CoT) prompting (Kojima et al., 2022) and include the sentence "Start your answer with lets analyze the image." into the prompt. If feasible, we proceed as before. If not, the LLM receives a list of all object parts rejected so far and is asked to propose another grasp. This loop repeats until either a feasible grasp part is found or there are no more suitable parts of the object left for grasping. In our implementation, the roles of the LLM and the VLM are carried out by the same model (GPT-4o).
To evaluate this approach, we collected a dataset of $18$ challenging scenarios, shown in Figs.˜9 and 8, where the initial grasp suggestion is not feasible and requires reasoning to find the right grasp part. We defined the ground truth manually to evaluate the results. We ran the algorithm $5$ times to obtain an average success rate of $0.91$ over all scenes. The qualitative results are shown for a part of the dataset in Fig.˜8. Our algorithm can take into account different criteria, e.g., the risk due to the proximity of the lit candle, potential mess due to the overfilled cup, or occlusions by nearby objects. We further analyze the effects of VLM choice and CoT in the following section.
@@ 4.6 Ablation Study
In this section, we evaluate the influence of different algorithm components on the above results, specifically the choice of the LLM and VLM and the effect of CoT on the feasibility feedback. First, we consider the pipeline without feedback as described in Section˜3.1. Here we consider GPT-3.5-turbo and GPT-4 as LLMs. We further investigate whether replacing the LLM with a VLM improves the performance. To that end, we deploy GPT-4o, GPT-4o-mini, and the open-source LLaVA-1.5 7B VLM and provide them with an image of the object to grasp.
Here, we run our pipeline only until detecting the grasping area in the image, without executing the grasp nor generating a grasping pose. Therefore, we use different metrics than in Section˜4.4. First, we count the exact matches between the LLM-generated and GT object part labels and report the average success rate. Second, we compute the Intersection-over-Union (IoU) for the proposed grasping regions and the ground truth. All algorithm versions are evaluated on data from Section˜4.1 and the results are summarized in Table˜2.
Both text-only GPT versions perform equally with an $0.82$ success rate and an IoU score of around $0.63$. The VLM variants perform slightly better with a success rate of $0.86$. We note that there is no difference between the flagship GPT-4o and the downsized GPT-4o-mini model. LLaVA performed significantly worse with $0.46$ success rate and $0.38$ IoU score. Analyzing the object parts suggested by LLaVA showed that the model was correct for objects possessing a handle, for other objects the answer was either wrong or referred to generic image locations, e.g., ’bottom’ or ’top’. A straightforward explanation could be simply the smaller model size. However, another reason might be that we used the same prompt for LLaVA as for the GPT models, and better results could be achieved with further prompt engineering specifically targeting LLaVA.
Table 2: Ablation of LLMs and VLMs in the main Lan-grasp pipeline.
Method Text only Text + Image
@@
For the feasibility feedback algorithm, we compare GPT-4o and GPT-4o-mini. Further, we experiment with two prompt variants, the first with CoT (zero-shot-CoT) and the second without (zero-shot). In the latter, we omit the sentence "Start your answer with lets analyze the image." from the prompt. The experiments were performed in the same fashion as in Section˜4.5 and the results are reported in Table˜3. First, we consider the CoT variants. With $0.63$ success rate, GPT-4o-mini performed significantly worse than the larger GPT-4o ($0.91$), which indicates that model size is an important factor for complex reasoning. Without CoT the performance dropped to $0.65$ for GPT-4o and $0.53$ for GPT-4o-mini. That result demonstrates that our algorithm, in fact, benefits from the CoT approach.
@@
Success rate 0.91 0.65 0.63 0.53
4.7 Object Placement
We carry out the object placement experiments on a Franka Research 3 robot platform equipped with a parallel gripper and custom gripper fingers. The Robot Control Stack framework (RCS) (Jülg et al., 2025) is used for controlling the robot. The scene is captured by a static RGB camera with the image plane facing straight towards the experimentation table plate. GroundingDINO (Liu et al., 2024) and SAM2 (Ravi et al., 2025) are used for object segmentation and ChatGPT 4.5 is deployed as the VLM.
Figure 10: Initial object configuration examples from our placement setup: for each evaluated object, we show the RGB scene at the beginning of a trial (object held in the gripper) and the corresponding SAM 3D reconstructed mesh, aligned with the object’s observed pose.
Figure 11: Example sequence from a single placing trial: (1) initial camera image with the grasped object, (2) the estimated pose from the first alignment cycle is applied, (3) a 90° yaw rotation reorients the grasp to prepare for the second cycle, (4) the estimated pose from the second alignment cycle is applied, (5) the object is placed on the surface, and (6) the gripper opens to release the object.
4.7.1 Placement Experiments
We evaluate the placement method on six objects that have a clear $up$-orientation, shown in Fig.˜10. For each object, we tested five initial object rotations. Using the robot world coordinate frame shown in Fig.˜3, the initial roll-pitch orientations (in degrees) are (0,30)(0,-30), $(30,0)$, (30,0)(-30,0), (30,30)(30,-30), and (30,30)(-30,-30). We first position the gripper in a neutral position, place the object upright between the gripper fingers, and finally rotate the arm to one of the above configurations. We thus consider these roll and pitch angles as ground truth values for placement rotations.
For each object and initial configuration, we execute the method described in Section˜3.3. First, we retrieve the 3D model of the object by prompting SAM 3D with a descriptive prompt, such as cup, potted plant, or pink wine glass. We found that segmentation can sometimes fail when too little of the object is visible, either due to rotation or occlusion by the gripper. In these cases, we made slight adjustments to the grasp and repeated the experiment. We only report a failure if the segmentation of an object fails consistently, as is the case with the transparent wine glass.
After retrieving the 3D model, we perform one alignment cycle, rotate the gripper by $90^{\circ}$ to reorient the grasp, perform the second cycle, and place the object onto the table by lowering it until the robot detects a contact and stops. Finally, we open the gripper. We count a trial as successful if the object remains standing upright after being released from the gripper.
Table 4: Main results of placement experiments. The first row shows the placing success rate per object over five initial object configurations. The second and third rows report the average alignment error per object after one alignment cycle, for roll and pitch respectively.
Object potted plant bottle flower vase pink wine glass cup wine glass mean
Success rate 1.0 1.0 0.8 1.0 1.0 0.4 0.87
Roll error $15.03^{\circ}$ $12.09^{\circ}$ $3.56^{\circ}$ $4.50^{\circ}$ $9.12^{\circ}$ $11.05^{\circ}$ $\textbf{9.30}^{\circ}$
Pitch error $3.53^{\circ}$ $1.51^{\circ}$ $35^{\circ} 2.2$$22^{\circ}$ $3.05^{\circ}$ $1.56^{\circ}$ $\textbf{2.43}^{\circ}$
4.7.2 Placement Results
The per-object success rates are summarized in Table˜4. For all but two objects, all five trials were successful. For vase, one trial failed because the VLM selected the wrong $up$-orientation. In case of wine glass, two trials failed due to failed object segmentation and one trial due to an imprecise object pose estimation by SAM 3D. Averaged on all objects and trials, our method achieves a success rate of $0.87$. As shown in Table˜4, our method achieves high success rates for most objects. However, the wine glass proved to be the most challenging case due to its transparency, which provides limited visual cues.
We also analyze the orientation estimation error for roll and pitch. In Table˜4, we report the average roll and pitch errors per object after the first alignment cycle, where we leave out the trials failed due to VLM or object segmentation. Averaged on all objects, SAM 3D achieves a pitch error of $2.43^{\circ}$ but a significantly larger roll error of $9.3^{\circ}$. Intuitively, the roll is more difficult to estimate, as less of the object is visible in the image.
4.7.3 Object Frame Alignment
Our placement method assumes that SAM 3D creates the local object frame such that the $up$-direction coincides with one coordinate axis. This is not a feature reported in the original paper but an observation we made by evaluating over $30$ objects with diverse classes and shapes for which a clear $up$-direction can be defined. For all tested object, this alignment property was present. Some of the object meshes along with the object frame are depicted in Fig.˜12. We hypothesize that the model mimics the object poses seen in training data, which has been composed from stock 3D object libraries and real-world scenes. In both cases, it is likely that the objects would be oriented in a natural way, i.e., upright. Fig.˜12 also compares the SAM 3D object frame with coordinate frames derived from the aligned bounding box and principal components of the mesh vertices. There, the first four objects have a regular shape and exhibit rotational symmetry, thus the frames estimated through bounding boxes and PCA mostly coincide with that of SAM 3D. However, for objects with irregular shape, shown in the last two rows, the bounding box and PCA frames are not aligned with the $up$-direction. These examples demonstrate that, in general, the $up$-direction cannot be determined by geometric analysis alone and requires semantic understanding.
Input SAM 3D Bounding Box PCA
Refer to caption Refer to caption Refer to caption Refer to caption
$0.54^{\circ}$ $0.1^{\circ}$
Refer to caption Refer to caption Refer to caption Refer to caption
$77^{\circ} 0.0$$04^{\circ}$
Refer to caption Refer to caption Refer to caption Refer to caption
$1.76^{\circ}$ $1.62^{\circ}$
Refer to caption Refer to caption Refer to caption Refer to caption
$10.74^{\circ}$ $2.14^{\circ}$
Refer to caption Refer to caption Refer to caption Refer to caption
$31.68^{\circ}$ $29.16^{\circ}$
Refer to caption Refer to caption Refer to caption Refer to caption
$44.09^{\circ}$ $32.45^{\circ}$
Figure 12: Comparison of object coordinate frames obtained by different methods. Columns from left to right: Input image, object mesh and frame by SAM 3D, bounding box frame, and PCA frame. The numbers in the last two columns are angular deviations from the SAM 3D frame, measured between the $y$-axes.